### v2

# xRFM: Accurate, scalable, and interpretable feature learning models for tabular data

Abstract
Inference from tabular data, collections of continuous and categorical variables organized into matrices, is a foundation for modern technology and science. Yet, in contrast to the dramatic developments in the rest of AI, the best practice for these predictive tasks has been relatively unchanged and is still primarily based on variations of Gradient Boosted Decision Trees (GBDTs). Very recently, there has been renewed interest in developing state-of-the-art methods for tabular data based on recent developments in neural networks and feature learning methods. In this work, we introduce xRFM, an algorithm that combines feature learning kernel machines with a tree structure to both adapt to the local structure of the data and scale to essentially unlimited amounts of training data.
On the TALENT benchmark, we show that compared to $31$ other methods, including recently introduced tabular foundation models (TabPFNv2) and GBDTs, xRFM achieves best performance across $100$ regression datasets and is competitive to the best methods across $200$ classification datasets outperforming GBDTs. Additionally, xRFM provides interpretability natively through the Average Gradient Outer Product. Code for xRFM (following a scikit-learn-style API) is available at: https://github.com/dmbeaglehole/xRFM.
1 Introduction
Tabular data – collections of continuous and categorical variables organized into matrices – underlies all aspects of modern commerce and science from airplane engines to biology labs to bagel shops. Yet, while methods for Machine Learning and AI in language and vision have seen unprecedented progress, the primary methodologies of prediction from tabular data have been relatively static, dominated by variations of Gradient Boosted Decision Trees (GBDTs), such as XGBoost [8]. Nevertheless, hundreds of tabular datasets have been assembled to form extensive regression and classification benchmarks [6, 18, 14, 38, 13], and, recently, there has been renewed interest in building state-of-the-art predictive models for tabular data [20, 21, 17]. Notably, given the remarkable effectiveness of large, “foundation” models for text, there has been much excitement in developing similar models on tabular data, and recent effort has led to the development of TabPFN-v2, a foundation model for tabular data appearing in Nature [20]. Yet, despite this progress, building scalable, effective, and interpretable machine learning models in this domain remains an open challenge.
In this work, we introduce xRFM, a tabular predictive model that combines recent advances in feature learning kernel machines with an adaptive tree structure, making it highly accurate, scalable, and interpretable. xRFM builds upon the Recursive Feature Machine (RFM) algorithm from [29], which enabled feature learning (a form of supervised dimensionality reduction) in general machine learning models. xRFM works as follows: Given training data, it first builds a binary tree structure to split data into subsets based on features relevant for prediction within each split. When splits reach a certain size (of $\leq C$ data samples), we train a leaf RFM (a hyper-parameter and compute optimized version of the original RFM).
There are two important consequences of using a tree structure in conjunction with leaf RFMs. The first is that it enables local feature learning, meaning that xRFM can learn different features for different subsets of the data. This property is crucial for tackling hierarchical structure in tabular data (for example, one set of features may be relevant when a specific feature takes on high value while another set of features is relevant when the same feature takes low values). The second consequence is that it enables xRFM to scale nearly linearly in the number of samples ($O(n\log n)$ given $n$ samples) during training and logarithmically ($O(\log n)$) during inference. As such, the inference time of xRFM is comparable to that of tree-based models.
In practice, we show xRFM has the best performance across 100 tabular regression tasks and is competitive with state-of-the-art on 200 tabular classification tasks from the TALENT benchmark [38]. In particular, xRFM outperforms 31 other methods on these regression tasks including the GBDT and neural network variations mentioned in the opening paragraph. On the TabArena-Lite benchmark [13], we show xRFM achieves one of the best tradeoffs (is on the empirical Pareto frontier) between performance and inference time among all methods in regression and is again competitive on classification. We show that xRFM achieves similar results on the largest datasets from the meta-test benchmark [21], where directly solving standard kernel machines becomes intractable on standard GPUs. An additional benefit of xRFM is that it natively provides interpretability by exposing features learned and used for prediction. In particular, each leaf RFM learns features through a mathematical object known as the Average Gradient Outer Product (AGOP), whose diagonal indicates coordinates relevant for prediction and whose top eigenvectors indicate directions in data most relevant for prediction. Using examples of xRFM trained on synthetic and real data, we show how AGOP matrices at each leaf RFM shed light on features relevant for prediction, including cases when leaf RFMs learn different features for different splits of data.
To summarize, xRFM combines feature learning kernel methods with adaptive data partitioning using the AGOP. It is a fast, effective, and interpretable model on tabular data at all scales.
2 Preliminaries
We begin with a review of kernel machines and kernel-RFM, which we use to build xRFM.
2.1 Kernel machines
A kernel machine is a non-parametric machine learning model [33, 3]. The idea behind kernel machines is that a nonlinear predictive model can be trained by first transforming input data with a fixed, nonlinear feature map and then performing linear regression on the transformed data. Kernel machines make this procedure computationally tractable even for infinite dimensional feature maps by using explicitly defined kernel functions (inner products of feature mapped data). We describe kernel machines in the context of supervised learning below. Let $X\in\mathbb{R}^{n\times d}$ denote training inputs with ${x^{(i)}}^{T}$ denoting the example in the $i$th row of $X$ for $i\in[n]$ and $y\in\mathbb{R}^{n\times c}$ denote training labels. Let $K:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ denote a kernel function (a positive semi-definite, symmetric function). Given a regularization parameter $\lambda\geq 0$, a kernel machine trained on the data $(X,y)$ is a predictor, $\widehat{f}:\mathbb{R}^{d}\to\mathbb{R}^{c}$, of the form:
$\widehat{f}(x)=K(x,X)\alpha\qquad;\qquad\alpha=(K(X,X)+\lambda I)^{-1}y~;$ (1)
where the notation $K(x,X)\in\mathbb{R}^{1\times n}$ denotes an $n$-dimensional row vector with $K(x,X)_{i}=K(x,x^{(i)})$ and $K(X,X)\in\mathbb{R}^{n\times n}$ denotes a matrix with $K(X,X)_{ij}=K(x^{(i)},x^{(j)})$. Examples of kernel functions used in practice are the Gaussian kernel ($K(x,z)=\exp(-\|x-z\|_{2}^{2})/L^{2}$) or the Laplace kernel ($K(x,z)=\exp(-\|x-z\|_{2})/L$). The advantage of kernel functions is that the predictor admits a closed form solution, which can be computationally efficient to fit on datasets under 100100k samples. The limitations of kernel machines are twofold: (1) the formulation in Eq. (1) does not scale well (it is super-quadratic in the number of samples) and (2) the choice of kernel is fixed independently of the input data. The former limitation is addressed by recent, pre-conditioned gradient descent methods for solving for $\alpha$ in Eq. (1) and methods that construct low-dimensional approximations of the kernel matrix. These include EigenPro [24, 25, 2], Falkon [31], and randomly pivoted Cholesky [9]. The latter limitation is addressed by procedures for adapting the kernel to the training data. We describe one such approach known as the kernel Recursive Feature Machine (kernel-RFM) below.
2.2 Recursive Feature Machines (RFMs)
The ability to learn task-relevant features from data is key to building effective predictors [10, 15, 1]. RFM, introduced in Radhakrishnan et al. [29], is an algorithm that enables feature learning in general machine learning models through a mathematical object known as the Average Gradient Outer Product (AGOP). Given a predictive model $\widehat{f}:\mathbb{R}^{d}\to\mathbb{R}$ and data $S=\{x^{(1)},\ldots,x^{(n)}\}\subset\mathbb{R}^{d}$, the AGOP is defined as
$\text{AGOP}(\widehat{f},S)=\frac{1}{n}\sum_{i=1}^{n}\nabla\widehat{f}(x^{(i)})\nabla\widehat{f}(x^{(i)})^{T}\in\mathbb{R}^{d\times d},$ (2)
where $\nabla\widehat{f}(x^{(i)})$ denotes the gradient of $\widehat{f}$ at the point $x^{(i)}$. The AGOP is an estimate of the (un-centered) covariance of the gradients of $\widehat{f}$ and intuitively captures the subspace along which the predictor highly varies [36, 22]. The RFM algorithm involves iterating between training a predictive model and using the AGOP of the trained model to select features and linearly transform input data. As such, feature learning through RFM can be viewed as a form of “supervised” PCA.
In the particular case of kernel machines, which have no native mechanism for feature learning, RFM enables feature learning by adapting the kernel to the data. We describe the kernel-RFM algorithm below. Following the notation in the previous section, let $(X,y)$ denote training inputs and labels, let $K$ denote the kernel function, and $\lambda$ denote the regularization parameter. Letting $M_{1}=I$ and $c>0$, kernel-RFM repeats the following two steps for $T$ iterations:
$\text{Step 1:}~~\widehat{f}_{t}(x)=K(M_{t}x,XM_{t})\alpha_{t}\qquad;\qquad\alpha_{t}=[K(XM_{t},XM_{t})+\lambda I]^{-1}y~,$
$\text{Step 2:}~~M_{t+1}=\left[\text{AGOP}(\widehat{f}_{t}(M_{t}x),X)\right]^{c}.$ (3)
Typical choices of $T$ and $c$ in practice are $T\leq 10$ and $c\in\{\frac{1}{4},\frac{1}{2}\}$ [29, 5, 4, 26]. (In practice, we return the $f_{t}$ with best validation performance rather than returning $f_{T}$.) When training labels, $y$, are multi-dimensional, we use the Jacobian of $\widehat{f}$ instead of the gradient in Step (2) (i.e., averaging the AGOP over output dimensions). Kernel-RFM is particularly effective at identifying low dimensional subspaces (or subsets of variables) relevant for prediction [30, 39], making it a useful interpretability tool, as we will show in Section 5.
3 xRFM algorithm overview
Figure 1: Overview of xRFM training and inference procedures. (A) xRFM is trained by splitting the data along the median projections (denoted $c_{1},c_{2}$) onto computed split directions (denoted $v_{1},v_{2}$). Data is split repeatedly into leaves, which contain at most $C$ training samples. Leaf RFMs are trained on the data at each leaf. (B) During inference, test data is routed to the appropriate leaf RFM based on split directions. The prediction is generated by the selected leaf RFM.
We now describe our algorithm xRFM, which consists of the following two components (Fig. 1):
  1. (1)
    An improved kernel-RFM, termed leaf RFM, that is trained on the subset of data;
  2. (2)
    A binary tree that splits the data into subsets, termed leaves, of a maximum size (leaf size) that is independent of the number of training samples. Leaf RFMs are then trained on the subset at each leaf. The splits stratify data based on features relevant for prediction.
Together these components enable xRFM to perform local feature learning (learning different features on different subpopulations) and achieve $O(n\log n)$ training time as well as $O(\log n)$ inference time given $n$ samples. We outline these two components in detail below.
3.1 Leaf RFM
Here, we describe the changes we made to the original kernel RFM algorithm from Radhakrishnan et al. [29] to produce leaf RFMs. The kernel RFM model was primarily built using kernels that were invariant to orthonormal transformations of data (those where $K(x,z)=K(Ux,Uz)$ for any orthogonal matrix $U$). Examples include the Gaussian kernel of the form $K(x,z)=\exp(-L\|x-z\|_{2}^{2})$ or the Laplace kernel of the form $K(x,z)=\exp(-L\|x-z\|_{2})$. Such invariance does not effectively leverage a special property of tabular data – the fact that each coordinate can be independently meaningful. This property has been hypothesized as an explanation for the superior performance of tree-based models, which compose functions of individual coordinates, over traditional neural networks on tabular datasets [18]. To account for this special structure, we introduce the following modifications to kernel-RFM:
Figure 2: Training xRFM on synthetic data where splitting on the top AGOP direction enables xRFM to learn locally relevant features.
  1. (1)
    We tune over a more general class of kernels $K_{p,q}(x,x^{\prime})=\exp(-\|x-x^{\prime}\|_{p}^{q}/L^{q})$ that are positive definite for $0<q\leq p\leq 2$ [32, Theorems 1, 5].
  2. (2)
    We tune over using the full AGOP and just the diagonal of the AGOP. The latter is known to be a theoretically grounded approach for coordinate selection [39], and introduces an axis-aligned bias that has been observed to match the structure of tabular data [18].
In addition to the above changes, we also implemented optimizations to speed up computations for categorical variables and an adaptive approach for tuning bandwidth separately for data on each leaf. Additional details regarding these changes and hyperparameter search spaces for leaf RFMs on TALENT, TabArena-Lite, and meta-test are provided in Appendix A.
3.2 Tree-based data partitioning
We now explain how xRFM builds a binary tree structure to partition data into leaves on which leaf RFMs are trained. First, given a dataset $S$ with $n$ samples, we subsample $m$ points and train a leaf RFM (referred to as a split model) for one iteration on this subsample. The split model serves to learn a direction that can be used to partition the data into two subsets. To this end, we extract the top eigenvector, $v$, of the AGOP from the split model and create two subsets of the $n$ datapoints: $S_{1}=\{x\in S~;~v^{T}x>\text{Median}(v^{T}z~\text{for $z\in S)$}\}$ and $S_{2}=\{x\in S~;~v^{T}x\leq\text{Median}(v^{T}z~\text{for $z\in S)$}\}$. We repeat this procedure on $S_{1}$ and $S_{2}$ and their corresponding children until all leaves are less than a maximum leaf size $C$. We finally train a leaf RFM on each of these leaves. This procedure is illustrated in Fig. 1A and detailed in the ‘TreePartition’ procedure of Algorithm A.2.
Note that our splitting procedure results in leaves of equal sizes (because of the split based on median), unlike tree-based predictive models, which can result in uneven splits. We typically split large datasets until leaves contain at most 60,000 samples. Leaf RFM hyperparameters are tuned using only on the validation data that falls into the leaf (as determined by splits down the binary tree). In the case that validation data at a given leaf has too few samples, we further hold out a subset of the training data at the leaf to use for validation.
Note that it is possible to have split data based on unsupervised splitting procedures (e.g., those based on random projections or using top eigenvectors of PCA) [11]. A key advantage of our split approach is that it groups together data points based on the features most relevant for prediction, as is captured by the top eigenvector of the AGOP. We found that splitting in this manner outperformed splits based on unsupervised approaches, including those described above. Prior works have also split data using the random forest procedure. For example, Hollmann et al. [20] train TabPFN-v2 on the samples routed to a given leaf of a tree in a random forest. The procedure in xRFM differs as it stratifies samples by projection onto a direction rather than individual coordinates and produces only one balanced tree instead of a forest.
Figure 3: Performance and runtime of xRFM on the TALENT (Plots A-C) and TabArena-Lite benchmarks (Plots D-F). The y-axes of plots A-C are the shifted geometric mean of the error across all datasets in that category, while the x-axes are the average over all datasets of the training plus inference time per 1000 samples for just the best hyperparameter configuration (meaning if a dataset has $n$ samples, we compute the training and inference time on the $n$ samples divide the total time by $n/1000$). The y-axes in plots D-F are Elo, the main metric used in TabArena, and reflect the relative win-rate of each method, while the x-axes are the median inference time per 1000 total samples. The TabArena-Lite plots additionally show the default methods and the compute Pareto tradeoff curve for inference time versus Elo. The TALENT plots are (A) nRMSE over 100 regression datasets, (B) classification error over 80 multi-class datasets, (C) classification error over 120 binary classification datasets. The TabArena-Lite plots are (D) regression, (E) multi-class, and (F) binary datasets.
Advantages of tree splits over traditional methods for scaling kernels.
Because of tree-based partitioning, xRFM is able to scale log-linearly in the number of samples during training [^1] The complexity for the applications of RFM itself scales only linearly, but splitting the data with the median at every tree node increases the complexity to $O(n\log n)$. and logarithmically in the number of samples during inference. While this is a computationally appealing aspect in practice, we note that there are a number of existing methods for scaling kernel machines. Examples include the Nyström approximation [37] and fast preconditioner methods for solving kernel regression [24, 31, 25, 2].
A major advantage of xRFM over these other methods and kernel RFM itself is the ability for tree-based partitioning to learn features local to subsets of data. Here, in Fig. 2, we provide an example clearly demonstrating this benefit. In this example, the function $f:\mathbb{R}^{d}\to\mathbb{R}$ depends on coordinates $x_{1},x_{3},x_{5}$ if $x_{0}>0$ and depends on coordinates $x_{9},x_{11},x_{13}$ if $x_{0}\leq 0$. xRFM learns $x_{0}$ as the initial split direction (through the AGOP of the split model trained on a random subset), then learns the relevant features for each case in separate leaf RFM models. In contrast, if running kernel RFM from Radhakrishnan et al. [29], the model would simply return that $\{x_{0},x_{1},x_{3},x_{5},x_{9},x_{11},x_{13}\}$ are all relevant features without decomposing into the two cases.
4 xRFM performance
[…]
Figure 4: Comparisons of xRFM with (A) kernel ridge regression and (B) the original RFM [30] on TALENT regression datasets (metric is normalized $R^{2}$, see Appendix A). Each point is a dataset.
Performance evaluation measures.
When measuring performance on these datasets, we use Root Mean Square Error (RMSE) for regression tasks (as is used in TALENT), and we use classification error ($1$ - classification accuracy) for classification tasks. For TabArena-Lite, binary classification performance for individual datasets are measured by AUROC, while multi-class datasets are evaluted with log-loss. To measure performance on aggregate over TALENT, we consider the following aggregation metrics on individual dataset performance measures: Shifted Geometric Mean, Arithmetic Mean, and Normalized Arithmetic Mean. TabArena-Lite uses Elo [12], a rating system that calculates the relative skill levels of each method based on the probability of winning when pairs of methods are compared on individual datasets (see Erickson et al. [13] for how Elo is estimated in this benchmark). All other metrics are defined in Appendix A.
Results on the TALENT and TabArena-Lite benchmarks.
The TALENT benchmark consists of $300$ total datasets comparing $31$ different supervised learning algorithms [38]. Among these $31$ algorithms are strong, widely used predictive models including Gradient Boosted Decision Tree (GBDT) variants (like XGBoost, CatBoost, LightGBM), hyper-parameter optimized neural networks (RealMLP), and recent transformer-based foundation models (TabPFN-v2). The TALENT benchmark consists of $100$ regression tasks, $120$ binary classification tasks, and $80$ multiclass classification tasks. The number of training samples per task varies between $500$ and $100,000$ and each task contains at least $5$ input variables. TabArena-Lite contains $51$ total datasets ($13$ regression and $38$ classification) comparing 15 different supervised learning algorithms including many of those in TALENT. The datasets in TabArena-Lite vary in sample size from $700$ to $150,000$.
[…]
Fig. 5 shows that the overall training+inference cost of xRFM initially scales super-linearly. However, xRFM is still reasonably fast at 60K samples, after which the log-linear scaling thanks to the tree-based data partitioning kicks in (Fig. B.1). Notably, xRFM is extremely fast on small datasets: it is at least two orders of magnitude faster than other methods when the dataset contains fewer than 3000 samples (we omit TabPFN-v2 in our comparison as it could not be run on all datasets due to restrictions on the allowable sample size, number of features, and output dimensionality).
Results on large datasets from the meta-test benchmark.
To analyze xRFM performance on large datasets beyond those in TALENT and TabArena, we consider the $17$ largest datasets in the meta-test benchmark from [21] (Table B.6).
We compare the performance of xRFM to other models with results reported in the literature on these $17$ datasets (Fig. 6). These other models include GBDTs (XGBoost, LightGBM, CatBoost), and neural networks (ResNet, MLP, MLP-PLR [16], RealMLP). We report the percentage improvement of each method over MLPs, following the procedure in Gorishniy et al. [17], Qu et al. [28]. Similar to the results on TALENT, xRFM is best on regression tasks, and second best on classification tasks. All results are presented in Tables B.6B.7. For meta-test, we search over slightly larger regularization parameters and slightly smaller bandwidth parameters to adapt to the much larger sample sizes of these datasets.
Figure 5: Total training and inference time for the best hyperparameter configuration as a function of the number of samples (training+validation+testing) across the TALENT benchmark. Curves indicate piece-wise linear fit to measures on each dataset (shown as points). (A) Results across 100 regression tasks. (B) Results across 80 multi-class classification tasks. (C) Results across 120 binary classification tasks.
Figure 6: Performance comparison across the $17$ large datasets from meta-test (70,000-500,000 samples). (A, B) Percentage improvement over MLP error on (A) regression and (B) classification datasets. Average percentage improvement is denoted as a diamond. Points denote individual dataset results (points are jittered for visibility).
5 Features learned by xRFM on tabular data
An advantage of xRFM is that it immediately provides a means of identifying features relevant for prediction (without stacking on any additional interpretability methods). Namely, we can extract learned AGOPs of leaf RFMs and visualize the features they select.
We use two approaches for identifying features selected by the AGOP. The first is to identify the elements along the diagonal of the AGOP with highest magnitude. By definition, these entries indicate how much a leaf RFM’s predictions vary when perturbing a given coordinate. As such, they are a natural measure of feature importance. The second approach is to examine the loadings onto the top eigenvectors of the AGOP. This approach allows us to identify joint effects of feature perturbations on the prediction. Namely, if coordinate $i$ of the top eigenvector is positive and coordinate $j$ is negative, then increasing one of these features and decreasing the other changes the prediction.
We use both approaches above to understand what features are learned by xRFM on tabular data from scikit-learn and meta-test (Fig. 7). As examples, we study the AGOPs for four such datasets: (1) California housing - a regression task for predicting the average price of a house, (2) Covertype - a multiclass classification task for identifying the dominant tree species in a given location, (3) NYC Taxi Tipping - a regression task for predicting the dollar tip amount in a taxi ride, and (4) Breast cancer - a classification task for identifying malignancy from features of biopsy images.
As the California housing and breast cancer datasets contain fewer than 5050k samples (the parameter we used for leaf size), we visualize a single AGOP. For covertype, we visualize one of the AGOPs from a leaf RFM (for this dataset, the leaf RFM AGOPs generally indicate the same pattern), and for taxi tipping we visualize several leaf RFM AGOPs. Upon visualization, it is apparent that AGOPs indicate low rank structure: either they highlight a subset of coordinates relevant for prediction (Fig. 7A, B) or they have a decay in the eigenvalue spectrum (Fig. 7C).
In Fig. 7A and B, we list feature names for the features with highest diagonal AGOP entries (darker shades of red indicate higher values). In all cases, we find that AGOP identifies sensible features for prediction. For the California housing dataset, xRFM identifies longitude, or east-west location, as the most important feature for predicting the average price of a house. Given that beach fronts (and major cities) in California are typically located to the west, the importance of house longitude is consistent with the hypothesis that homes closer to the beach are more expensive on average. For the Covertype prediction dataset, the example AGOP from a leaf RFM shows that elevation, distance to roadways, and distance to firepoints are the most important features. This finding is consistent with the hypothesis that elevations with different climates and the existence of fires / roadways can significantly affect viability of different tree species. For the taxi tipping dataset, we observe that leaf RFMs identify varying local features. For example, one leaf RFM (denoted Taxi Tipping 1 in Fig. 7B) selected pickup location as an important feature, while this feature was less important for a different leaf RFM (Taxi Tipping 3). Furthermore, fare code and the MTA tax have varying feature relationships at each leaf: in leaf RFM 1 and 2, the fare code and MTA tax has neutral or synergistic effects on the prediction value, while for leaf RFM 3, increasing fare code has the opposite effect on prediction as increasing MTA tax (shown as a blue square in Fig. 7B).
Figure 7: Interpreting xRFM through the AGOP of its constituent Leaf RFM models. (A) Examining the most important features for xRFM trained on California Housing (price prediction) and Covertype (dominant tree species prediction) datasets, based on the magnitude of diagonal entries. (B) Examining the features identified across three different Leaf RFM models for the NYC Taxi Tipping dataset. (C) Examining features learned for Breast Cancer detection from processed FNA imaging. The spectrum of this AGOP is plotted and the top eigenvector is shown in a bar plot. The most positive and negative entries of this eigenvector are boxed.
For the breast cancer dataset, the most important feature found by xRFM is the mean number of concave points in the biopsy image, which has been shown to be a significant indicator of malignancy [27]. When examining the top eigenvector of the AGOP, we find that standard error in the compactness measurements of cell nuclei is the feature with highest importance and has an opposite effect on malignancy as concavity (Fig. 7C). Our method suggests that benign cells have less uniform compactness than malignant cells.
6 Discussion
Tabular data has historically provided a proving ground for the development of novel predictive models. However, even as the last years have seen spectacular improvements in language processing and computer vision, the progress in tabular data has been far more modest. While this may be changing with the developments such as hyperparameter-optimized neural networks (e.g., RealMLP [21]), and very recently, pre-trained transformer-based tabular networks, such as TabPFN-v2 [28], there is still large scope for improved tabular data prediction.
Kernel machines provide a powerful conceptually and computationally elegant approach for predictive modeling – one simply transforms data with a nonlinear feature map and then performs linear regression. Yet, historically, these methods have not out-performed Gradient Boosted Decision Trees (GBDTs), while also being significantly more difficult to scale. There are two key limitations of standard kernel machines: (1) standard kernel choices lack adaptivity to features relevant for prediction, and (1) naive implementations scale super-quadratically in the number of training samples, making them difficult to train on larger datasets (past 7070k training samples). Significant progress in overcoming limitation (1) was made in recent work [29], which introduced Recursive Feature Machines (RFM) to enable feature learning in a range of models, including kernel machines. In fact, feature learning through RFM can be provably more sample efficient than standard kernels [39]. There has also been significant work addressing the limitation (2) including [37, 25, 31], to reference a few.
In this work, we have taken a different, tree-based approach to overcome these limitations. Our approach enables scalability to large datasets (log-linear training time and logarithmic inference time) while also leveraging the strength of RFM in identifying and exploiting local features. Indeed, the ability of xRFM to learn different features for different subpopulations of data through AGOPs at leaves is useful for characterizing heterogeneity across large datasets. Building on this core novelty and other algorithmic optimizations, we discuss a number of potential improvements and follow ups below.
A direction for future improvement of xRFM is to explore the use of iterative kernel solvers to train leaf RFMs. By using early stopping, these methods could obviate the need for tuning ridge regularization. The use of the adaptive binary tree structure to split data also opens up new avenues for exploration. For instance, it would be interesting to characterize the tradeoff between leaf size and performance and to develop new methods for determining when to stop splitting based statistics of leaf data. It might also be useful to consider alternative bandwidth selection methods for each leaf, for example by optimizing for bandwidths directly as in [19]. Lastly, it would be important to understand the effectiveness of AGOP as an interpretability mechanism, particularly in comparison to other widely used methods such as Shapley values [23], gradient-based importance scores [34, 35], and feature importance scores for tree-based models [7].
We also note that predictive modeling on tabular data is a highly active research area. There are continually new benchmarks and baselines, tuning, and cleaning methods to consider (e.g., using inner cross-validation, using metrics beyond accuracy for evaluating classification models, and studying weighted ensembles of models [13]). It is important to benchmark xRFM as such new settings arise.
Overall, we have shown that xRFM is an effective algorithm for inference from tabular data that scales to essentially unlimited data sizes and achieves performance exceeding or comparable to the current state-of-the-art. It combines the advantages of tree-based methods with the power and elegance of feature-learning kernel machines. We envision that xRFM will be used for both high-performing predictive modeling and uncovering heterogeneous structure in large-scale tabular data.
Acknowledgments
We would like to thank Jingang Qu, Ingo Steinwart, Tizian Wenzel, and Gaël Varoquaux for relevant discussions. We thank Charles Durham for contributing the KerMac library, which led to drastic speed up in kernel evaluations and gradient computations. We acknowledge support from the National Science Foundation (NSF) and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and #814639 as well as the TILOS institute (NSF CCF-2112665), the Office of Naval Research (ONR N000142412631) and DARPA (HR001125CD020). This work used the programs (1) XSEDE (Extreme science and engineering discovery environment) which is supported by NSF grant numbers ACI-1548562, and (2) ACCESS (Advanced cyberinfrastructure coordination ecosystem: services & support) which is supported by NSF grants numbers #2138259, #2138286, #2138307, #2137603, and #2138296. Specifically, we used the resources from SDSC Expanse GPU compute nodes, and NCSA Delta system, via allocations TG-CIS220009.
Appendix A Methods
A.1 Additional leaf-RFM modifications
Below, we provide further detail on the RFM modifications we made to implement leaf RFM.
  1. (1)
    We tune over a more general class of kernels $K_{p,q}(x,x^{\prime})=\exp(-\|x-x^{\prime}\|_{p}^{q}/L^{q})$ that are positive definite for $0<q\leq p\leq 2$ [32, Theorems 1, 5]. We typically search over $p\in\mathcal{U}(0.7,1.4)$ and $p\in\{q,2\}$ (Table A.1).
  2. (2)
    We tune over using the full AGOP and just the diagonal of the AGOP. The latter is known to be a theoretically grounded approach for coordinate selection [39], and introduces an axis-aligned bias that has been observed to match the structure of tabular data [18].
We also make the following changes to improve leaf RFM runtime on tabular data and to allow xRFM to adapt to the variance of input variables at each leaf:
  1. (3)
    We speed up computation for kernels with $q=1$ on categorical variables taking $c$ values by precomputing possible kernel entries as follows. We restrict the AGOP matrix to be block diagonal with a block for the $c\times c$ entries corresponding to the categorical variable. Letting $M_{c}\in\mathbb{R}^{c\times c}$ denote this block and $e_{i}\in\{0,1\}^{c}$ denote the one-hot embedding of the categorical variable when it takes on value $i$, we precompute $M_{c}^{1/2}(e_{i}-e_{j})$ for all $i,j\in[c]$.
  2. (4)
    We tune over whether or not to use adaptive bandwidth, which involves separately scaling $L$ at each leaf RFM by $\left(\text{Median}(\|x-x^{\prime}\|_{p}))\right)^{-1}$ for $x\neq x^{\prime}$ at every leaf independently. Adaptive scaling allows xRFM to adapt to the variance of covariates at different leaves.
A.2 Metrics
Following [21], we report the shifted geometric mean of the error, which is the geometric mean of the error after shifting by a small value to prevent over-sensitivity to datasets with small errors. This metric is defined as follows.
Definition 1.
For a given set of errors on a benchmark $\varepsilon_{1},\ldots,\varepsilon_{N}$, the shifted geometric mean ($\mathrm{SGM}_{\varepsilon}$) with parameter $\varepsilon$ takes value:
$\mathrm{SGM}_{\varepsilon}=\exp\left(\frac{1}{N}\sum_{i=1}^{N}\log(\varepsilon+\varepsilon_{i})\right)~.$
In this work, as in Holzmüller et al. [21], we use $\varepsilon=0.01$.
For regression tasks, we use normalized Root-mean-square-error (nRMSE), which is defined as follows.
Definition 2.
The normalized Root-mean-square-error is defined as,
$\text{nRMSE}\;=\;\sigma_{y}^{-1}\,\sqrt{\frac{1}{N}\sum_{i=1}^{N}\bigl(y_{i}-\widehat{y}_{i}\bigr)^{2}},\qquad\text{where}\;\sigma_{y}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}$
When we refer to other normalized metrics, such as those used in Table B.1, Table B.2, Table B.3, Table B.4, Table B.5, we are min-max normalizing errors across methods for each dataset (see below).
Definition 3.
The normalized error $\widetilde{E}_{j}$ for method $j$ on a given dataset, where the un-normalized error is $E_{j}$, has the form:
$\widetilde{E}_{j}\;=\;\frac{E_{j}-E_{\min}}{E_{\max}-E_{\min}},\qquad E_{\min}=\min_{k}E_{k},\quad E_{\max}=\max_{k}E_{k}~.$
A.3 Hyperparameters
Hyperparameter TALENT TabArena-Lite Meta-test
Bandwidth $\log\mathcal{U}(1,200)$ $\log\mathcal{U}(0.5,200)$ $\log\mathcal{U}(0.4,80)$
Bandwidth Mode {constant} {constant} {constant, adaptive}
Categorical Transformations {one_hot} {one_hot} {ordinal_encoding, one_hot}
Diagonal {False, True} {False, True} {False, True}
Early Stop Multiplier 1.1 1.1 1.1
Exponent $q$ $\mathcal{U}(0.7,1.4)$ $\mathcal{U}(0.7,1.4)$ $\mathcal{U}(0.7,1.3)$
[…]
$p$ (when kernel type is $K_{p,q}$) $\mathcal{U}(q,q+0.8(2-q))$ $\mathcal{U}(q,q+0.8(2-q))$ $\mathcal{U}(q,q+0.8(2-q))$
Normalization {standard} {standard} {standard}
Regularization $\log\mathcal{U}(10^{-6},1)$ $\log\mathcal{U}(10^{-6},10)$ $\log\mathcal{U}(10^{-5},50)$
Refill size ($N_{\mathrm{val}}$) 1500 1500 1500
Min. split subset size ($C$) 60,000 60,000 60,000
Table A.1: Search spaces for xRFM on the TALENT (Figure 3), TabArena-Lite (Figure 3), and Meta-test benchmarks (Figure 6).
Data pre-processing.
For the TALENT benchmark, each method’s hyperparameters are tuned after fixing a single data normalization and categorical encoding scheme. For xRFM, kernel ridge regression, and standard kernel RFM, we one-hot encode categorical features and z-score input coordinates separately. For meta-test, methods also tune over choice of ordinal or one-hot encoding of categorical variables. On meta-test, we also normalize data by z-scoring input coordinates separately prior to tuning xRFM parameters.
Details of various methods.
For scaling kernel ridge regression and the standard kernel RFM to large TALENT datasets, we used Eigenpro-2 (EP2) [25] that was initialized with the coefficients obtained from directly solving RFM on a random sample of 70,000 points. To tune the optimization hyperparameters for EP2, we tuned the model parameters for 100 trials on a random 70,000 sample subset, then for the final run, we used the tuned hyperparameters (and for kernel RFM the AGOP from the best iteration). For TabPFN-v2, we used the code provided from the benchmark github (https://github.com/LAMDA-Tabular/TALENT), which subsampled to 10,000 samples to avoid out-of-memory issues. The provided code did not apply TabPFN-v2 to datasets with more than 10 classes. For efficient $L_{p}^{q}$ kernel computations on GPU we used the KerMac library (https://github.com/Kernel-Machines/kermac).
A.4 Algorithmic details
Note on AGOP computation
For computing gradients of the predictor on training data (for AGOP computation), we omit the contribution to the gradient from the kernel evaluation between each training point and itself. This is because the kernel function is often not differentiable (e.g. Laplace kernels) when evaluated for two identical points.
Algorithm A.1 Leaf RFM
  • $x^{(1)},\ldots,x^{(n)}\in\mathbb{R}^{d},y\in\mathbb{R}^{n\times c}$ : Train data
  • $X_{\mathrm{val}}\in\mathbb{R}^{m\times d},y_{\mathrm{val}}\in\mathbb{R}^{m\times c}$ : Validation data
  • $K(\cdot,\cdot;L,p)$: Kernel parametrized by bandwidth $L\in\mathbb{R}^{+}$ and exponent $p\in(0,2]$
  • $\tau\in\mathbb{Z}^{+}$: Number of iterations
  • $\lambda\in\mathbb{R}^{+}$: Ridge parameter
  • $\varepsilon\in\mathbb{R}^{+}$: Stability parameter
  • use_diag: Boolean indicating whether to use diagonal of AGOP only
  • adapt_bandwidth: Boolean indicating whether to adapt bandwidth
$M_{0}\leftarrow I_{d\times d}$
$X=[x^{(1)},\ldots,x^{(n)}]^{\top}\in\mathbb{R}^{n\times d}$
if adapt_bandwidth then
  $L\leftarrow\mathrm{AdaptBandwidth}(L)$ $\triangleright$ Adapt bandwidth if enabled
end if
for $t=0,\ldots,\tau-1$ do
  if use_diag then
   $X_{M}\leftarrow X\odot\text{diag}(M_{t})^{1/2}$
   Solve $\alpha_{t}$ such that $(K(X_{M},X_{M})+\lambda I)\alpha_{t}=y$
   $f^{(t)}(x)=K(x\odot\text{diag}(M_{t})^{1/2},X_{M})\alpha_{t}$ $\triangleright$ $\odot$ denotes element-wise multiplication
  else
   $X_{M}\leftarrow XM^{1/2}_{t}$
   Solve $\alpha_{t}$ such that $(K(X_{M},X_{M})+\lambda I)\alpha_{t}=y$
   $f^{(t)}(x)=K(M^{1/2}_{t}x,X_{M})\alpha_{t}$
  end if
  Compute $E_{t}\leftarrow\mathrm{Error}(f^{(t)},X_{\mathrm{val}},y_{\mathrm{val}})$ $\triangleright$ Validate model
  $M_{t+1}\leftarrow\frac{1}{n}\sum_{i=1}^{n}\nabla_{x}f^{(t)}(x^{(i)})\nabla_{x}f^{(t)}(x^{(i)})^{\top}\in\mathbb{R}^{d\times d}$ $\triangleright$ Feature matrix (AGOP) computation
  $M_{t+1}\leftarrow M_{t+1}/(\varepsilon+\max_{i,j}M_{t+1}[i,j])$ $\triangleright$ Normalize feature matrix
end for
$t^{*}\leftarrow\operatorname*{arg\,min}_{t}E_{t}$
return $\alpha_{t^{*}},M_{t^{*}}$ $\triangleright$ KRR coefficients: $\alpha_{t^{*}}$, feature matrix: $M_{t^{*}}$ from best iteration on val. set
Algorithm A.2 TreePartition
  • $\mathcal{D}=\{(x^{(1)},y^{(1)}),\ldots,(x^{(n)},y^{(n)})\}$ : Full dataset with $x^{(i)}\in\mathbb{R}^{d},y^{(i)}\in\mathbb{R}^{c}$
  • $K(\cdot,\cdot;L,p)$: Kernel parametrized by bandwidth $L\in\mathbb{R}^{+}$ and exponent $p\in(0,2]$
  • $N\in\mathbb{Z}^{+}$: Number of sample points for Leaf RFM
  • $L\in\mathbb{Z}^{+}$: Maximum leaf size
  • $\lambda\in\mathbb{R}^{+}$: Ridge parameter
function TreePartition($\mathcal{D}$)
  if $|\mathcal{D}|\leq L$ then
   return Leaf node with dataset $\mathcal{D}$
  end if
  Sample NN points $\mathcal{S}=\{(x^{(a_{1})},y^{(a_{1})}),\ldots,(x^{(a_{N})},y^{(a_{N})})\}$ from $\mathcal{D}$
  $X_{s}=[x^{(a_{1})},\ldots,x^{(a_{N})}]^{\top}\in\mathbb{R}^{N\times d}$
  $y_{s}=[y^{(a_{1})},\ldots,y^{(a_{N})}]^{\top}\in\mathbb{R}^{N\times c}$
  Solve $\alpha$ such that $(K(X_{s},X_{s})+\lambda I)\alpha=y$ $\triangleright$ Fit Leaf RFM on sampled data
  Define predictor $f(x)=K(x,X_{s})\alpha$
  Compute AGOP: $M\leftarrow\frac{1}{N}\sum_{i=1}^{N}\nabla_{x}f(x^{(a_{i})})\nabla_{x}f(x^{(a_{i})})^{\top}\in\mathbb{R}^{d\times d}$
  Extract top eigenvector $v_{1}$ of $M$ $\triangleright$ Principal direction
  Project all data points: $p^{(i)}\leftarrow v_{1}^{\top}x^{(i)}$ for $i=1,\ldots,|\mathcal{D}|$
  Compute median projection: $m\leftarrow\text{Median}(\{p^{(1)},\ldots,p^{(|\mathcal{D}|)}\})$
  Split dataset:
   $\mathcal{D}_{\text{left}}\leftarrow\{(x^{(i)},y^{(i)})\in\mathcal{D}:v_{1}^{\top}x^{(i)}\leq m\}$
   $\mathcal{D}_{\text{right}}\leftarrow\{(x^{(i)},y^{(i)})\in\mathcal{D}:v_{1}^{\top}x^{(i)}>m\}$
  $\text{left\_child}\leftarrow\text{TreePartition}(\mathcal{D}_{\text{left}})$ $\triangleright$ Recursive call
  $\text{right\_child}\leftarrow\text{TreePartition}(\mathcal{D}_{\text{right}})$ $\triangleright$ Recursive call
return Internal node with splitting vector $v_{1}$, threshold $m$, and children
end function
Algorithm A.3 Route (find the leaf that contains a point)
  • $\mathcal{T}$ : a (possibly trained) tree whose internal nodes store
    • splitting vector $v_{1}\in\mathbb{R}^{d}$
    • threshold $m\in\mathbb{R}$
  • $x\in\mathbb{R}^{d}$ : query point
function Route($x,\mathcal{T}$)
  $r\leftarrow\mathcal{T}.\text{root}$ $\triangleright$ Initialize current node $r$ from the tree
  while $r$ is an internal node do
   if $v_{1}^{\top}x\leq r.\text{threshold}$ then $\triangleright$ Check if projection is less than (median) threshold
     $r\leftarrow r.\text{left\_child}$
   else
     $r\leftarrow r.\text{right\_child}$
   end if
  end while
  return $r$ $\triangleright$ $r$ is now the leaf node that contains $x$
end function
Algorithm A.4 xRFM (training)
  • $\mathcal{D}_{\mathrm{train}},\;\mathcal{D}_{\mathrm{val}}$ : training and validation sets
  • $K(\cdot,\!\cdot;L,p)$ : kernel (bandwidth $L$, exponent $p$)
  • $\text{TreeHyp}=\{N,\,L_{\max},\,\lambda{\text{split}}\}$ : hyper-parameters for TreePartition
  • $\text{LeafHyp}=\{\tau,\,\lambda_{\text{leaf}},\,\varepsilon,\,$use_diag}\} : hyper-parameters for LeafRFM
function xRFM-fit($\mathcal{D}_{\mathrm{train}},\mathcal{D}_{\mathrm{val}}$)
  $\mathcal{T}\leftarrow$ TreePartition($\mathcal{D}_{\mathrm{train}}$, $K$, TreeHyp) $\triangleright$ Alg. A.2
[…]
   $\mathcal{D}_{\ell}^{\mathrm{train}}\leftarrow$ data stored in $\ell$
   $\mathcal{D}_{\ell}^{\mathrm{val}}\leftarrow\bigl\{(x,y)\!\in\!\mathcal{D}_{\mathrm{val}}:\textsc{Route}(x,\mathcal{T})=\ell\bigr\}$
   $(\alpha_{\ell},M_{\ell})\leftarrow$ LeafRFM($\mathcal{D}_{\ell}^{\mathrm{train}},\mathcal{D}_{\ell}^{\mathrm{val}}$, $K$, LeafHyp) $\triangleright$ Alg. A.1
   Define predictor $f_{\ell}(x)$ according to $(\alpha_{\ell},M_{\ell})$ and store it in $\ell$
  end for
  return $\mathcal{T}$ $\triangleright$ Tree whose leaves now carry trained predictors
end function
Algorithm A.5 xRFM (inference)
  • $\mathcal{T}$ : trained tree returned by xRFM-fit (Alg. A.4)
  • $x\in\mathbb{R}^{d}$ : test point
function xRFM-predict($\mathcal{T},x$)
  $\ell\leftarrow\textsc{Route}(x,\,\mathcal{T})$ $\triangleright$ Alg. A.3
  $\widehat{y}\leftarrow f_{\ell}(x)$ $\triangleright$ Leaf predictor stored in $\ell$
  return $\widehat{y}$
end function
Appendix B All results
Method Rank Score Norm. Score Top-1 (%) Top-3 (%) Top-5 (%) Top-8 (%) $\mathrm{SGM}_{\varepsilon}$
xRFM 4.70 0.311 0.036 20.0 56.0 69.0 84.0 0.311
TabPFN-v2 6.55 0.323 0.067 20.0 45.0 57.0 72.0 0.323
CatBoost 7.79 0.336 0.053 7.00 24.0 41.0 68.0 0.336
RealMLP 8.20 0.329 0.076 8.00 21.0 37.0 60.0 0.329
ModernNCA 9.39 0.365 0.097 14.0 30.0 38.0 57.0 0.365
LightGBM 9.91 0.353 0.062 6.00 16.0 34.0 52.0 0.353
XGBoost 10.6 0.355 0.071 2.00 12.0 32.0 47.0 0.355
TabR 10.7 0.365 0.133 5.00 20.0 34.0 46.0 0.365
FTT 12.0 0.388 0.149 3.00 11.0 17.0 33.0 0.388
Laplace KRR 12.3 0.373 0.102 3.00 11.0 21.0 35.0 0.373
MLP-PLR 12.7 0.381 0.142 0.0 8.00 15.0 34.0 0.381
Excelformer 12.7 0.399 0.157 0.0 2.00 9.00 28.0 0.399
PTARL 13.1 0.390 0.114 2.00 6.00 9.00 20.0 0.390
RandomForest 13.3 0.389 0.088 5.00 12.0 22.0 35.0 0.389
AutoInt 14.3 0.399 0.158 1.00 2.00 6.00 12.0 0.399
Node 14.4 0.451 0.195 0.0 6.00 18.0 31.0 0.451
MLP 15.2 0.418 0.167 1.00 3.00 4.00 11.0 0.418
DCN2 15.3 0.473 0.227 0.0 2.00 10.0 20.0 0.473
ResNet 15.6 0.426 0.175 0.0 4.00 8.00 15.0 0.426
Tangos 16.3 0.432 0.173 0.0 1.00 4.00 9.00 0.432
SNN 17.2 0.419 0.174 0.0 2.00 4.00 8.00 0.419
kNN 19.1 0.459 0.190 2.00 4.00 4.00 11.0 0.459
TabNet 21.5 0.475 0.227 0.0 0.0 0.0 0.0 0.475
GrowNet 23.0 0.619 0.323 0.0 0.0 0.0 0.0 0.619
SVM 23.2 0.612 0.346 0.0 0.0 1.00 1.00 0.612
TabTransformer 26.4 1.07 0.750 1.00 1.00 2.00 3.00 1.07
SwitchTab 26.9 1.29 0.745 0.0 0.0 0.0 1.00 1.29
Danets 27.2 1.12 0.830 0.0 1.00 1.00 3.00 1.12
Table B.1: Full TALENT Regression results across 100 datasets. Rank is the average rank among the ordered methods over all datasets. Score is the metric we use to compare methods in Figure 3, in this case $SGM_{\epsilon}$. Normalized score is the arithmetic mean of the normalized nRMSE. Top-X (%) is the percentage of datasets for which that method is in the top X ranks. The final column is the shifted geometric mean error ($\mathrm{SGM}_{\varepsilon}$).
Method Rank Score Norm. Score Top-1 (%) Top-3 (%) Top-5 (%) Top-8 (%) $\mathrm{SGM}_{\varepsilon}$
TabPFN-v2 7.15 0.823 0.935 25.0 51.5 63.2 70.6 0.108
xRFM 7.60 0.825 0.933 11.8 29.4 48.5 66.2 0.107
RealMLP 7.64 0.823 0.918 10.3 20.6 48.5 66.2 0.107
TabR 7.98 0.828 0.932 8.82 27.9 41.2 60.3 0.106
ModernNCA 8.88 0.825 0.912 10.3 30.9 44.1 61.8 0.106
CatBoost 10.3 0.819 0.905 1.47 19.1 36.8 51.5 0.117
LightGBM 11.7 0.813 0.883 5.88 22.1 33.8 44.1 0.125
XGBoost 12.1 0.815 0.886 4.41 19.1 26.5 41.2 0.124
ResNet 12.3 0.807 0.879 0.0 4.41 16.2 36.8 0.127
FTT 12.8 0.810 0.868 0.0 5.88 10.3 29.4 0.125
MLP-PLR 13.2 0.815 0.881 0.0 7.35 11.8 25.0 0.119
Laplace KRR 13.5 0.804 0.864 5.88 16.2 20.6 33.8 0.133
DCN2 14.0 0.806 0.865 1.47 4.41 10.3 26.5 0.127
MLP 14.2 0.805 0.864 0.0 4.41 8.82 19.1 0.131
SNN 14.9 0.805 0.858 0.0 0.0 4.41 8.82 0.129
AutoInt 15.7 0.804 0.858 0.0 0.0 1.47 7.35 0.130
RandomForest 15.7 0.797 0.831 5.88 8.82 13.2 27.9 0.137
Excelformer 16.1 0.805 0.853 0.0 0.0 4.41 17.6 0.127
Tangos 16.5 0.797 0.844 0.0 1.47 4.41 13.2 0.133
TabCaps 16.5 0.797 0.837 0.0 1.47 5.88 13.2 0.134
Danets 17.7 0.795 0.829 0.0 1.47 2.94 5.88 0.137
PTARL 19.1 0.796 0.808 0.0 1.47 1.47 4.41 0.139
kNN 20.1 0.786 0.772 2.94 7.35 10.3 14.7 0.152
LogReg 20.2 0.767 0.755 1.47 4.41 7.35 13.2 0.162
TabTransformer 21.5 0.771 0.739 0.0 1.47 1.47 2.94 0.155
Node 22.0 0.767 0.742 0.0 4.41 7.35 10.3 0.173
SVM 23.4 0.750 0.697 2.94 5.88 5.88 8.82 0.182
GrowNet 24.4 0.691 0.596 0.0 0.0 2.94 5.88 0.216
SwitchTab 25.0 0.748 0.674 0.0 1.47 1.47 2.94 0.187
TabNet 26.0 0.753 0.615 0.0 0.0 0.0 0.0 0.183
NaiveBayes 29.4 0.609 0.327 0.0 0.0 0.0 0.0 0.321
NCM 30.5 0.627 0.278 0.0 0.0 1.47 1.47 0.326
Table B.2: Full TALENT Multiclass classification ($\leq 10$ classes) results. Rank is the average rank among the ordered methods over all datasets. (Normalized) Score is the metric we use to compare methods in Figure 3. In this case score is the mean classification accuracy (1 minus the error in Figure 3). Top-X (%) is the percentage of datasets for which that method is in the top X ranks. The final column is the shifted geometric mean error ($\mathrm{SGM}_{\varepsilon}$).
Method Rank Score Norm. Score Top-1 (%) Top-3 (%) Top-5 (%) Top-8 (%) $\mathrm{SGM}_{\varepsilon}$
RealMLP 3.75 0.730 0.964 8.33 50.0 91.7 91.7 0.164
xRFM 5.46 0.718 0.949 25.0 50.0 66.7 83.3 0.183
TabR 6.00 0.720 0.948 16.7 50.0 50.0 66.7 0.172
ResNet 6.33 0.720 0.949 0.0 16.7 50.0 83.3 0.178
ModernNCA 6.83 0.713 0.913 33.3 41.7 50.0 83.3 0.168
FTT 8.79 0.709 0.919 0.0 8.33 25.0 50.0 0.187
MLP-PLR 8.79 0.712 0.925 0.0 0.0 25.0 50.0 0.185
MLP 9.58 0.710 0.922 0.0 0.0 8.33 41.7 0.190
Laplace KRR 10.8 0.679 0.889 0.0 16.7 41.7 50.0 0.218
DCN2 11.2 0.707 0.915 0.0 8.33 8.33 8.33 0.193
AutoInt 12.2 0.701 0.899 0.0 0.0 0.0 16.7 0.194
SNN 12.2 0.704 0.906 0.0 0.0 0.0 33.3 0.197
CatBoost 12.8 0.679 0.842 8.33 25.0 33.3 33.3 0.218
Danets 15.9 0.687 0.882 0.0 0.0 0.0 0.0 0.215
Tangos 16.1 0.660 0.823 0.0 8.33 8.33 16.7 0.225
XGBoost 16.2 0.678 0.874 8.33 8.33 8.33 25.0 0.236
Excelformer 16.6 0.675 0.851 0.0 0.0 0.0 16.7 0.217
TabCaps 16.8 0.670 0.853 0.0 0.0 0.0 0.0 0.231
LightGBM 16.9 0.668 0.850 0.0 0.0 0.0 16.7 0.261
PTARL 18.7 0.656 0.823 0.0 0.0 0.0 0.0 0.233
kNN 20.2 0.637 0.800 0.0 0.0 0.0 8.33 0.273
RandomForest 20.7 0.620 0.782 0.0 8.33 8.33 8.33 0.311
TabTransformer 20.8 0.582 0.702 0.0 0.0 8.33 8.33 0.307
LogReg 22.0 0.538 0.622 0.0 0.0 0.0 0.0 0.338
SVM 25.1 0.486 0.523 0.0 8.33 8.33 8.33 0.385
SwitchTab 25.2 0.525 0.603 0.0 0.0 0.0 0.0 0.372
Node 26.2 0.536 0.611 0.0 0.0 0.0 0.0 0.402
GrowNet 26.2 0.402 0.477 0.0 0.0 0.0 0.0 0.565
TabNet 26.3 0.486 0.516 0.0 0.0 0.0 0.0 0.368
NaiveBayes 29.5 0.410 0.380 0.0 0.0 0.0 0.0 0.545
NCM 30.2 0.399 0.339 0.0 0.0 0.0 0.0 0.552
Table B.3: Full TALENT Multiclass classification (>10>10 classes) results. Rank is the average rank among the ordered methods over all datasets. (Normalized) Score is the metric we use to compare methods in Figure 3. In this case score is the mean classification accuracy (1 minus the error in Figure 3). Top-X (%) is the percentage of datasets for which that method is in the top X ranks. The final column is the shifted geometric mean error ($\mathrm{SGM}_{\varepsilon}$).
Method Rank Score Norm. Score Top-1 (%) Top-3 (%) Top-5 (%) Top-8 (%) $\mathrm{SGM}_{\varepsilon}$
xRFM 5.96 0.845 0.981 29.6 55.6 59.3 74.1 0.119
RealMLP 7.44 0.839 0.951 7.41 14.8 29.6 77.8 0.125
TabR 7.83 0.844 0.970 18.5 33.3 55.6 70.4 0.116
ModernNCA 9.69 0.843 0.963 3.70 29.6 40.7 59.3 0.120
CatBoost 10.2 0.827 0.891 7.41 25.9 33.3 55.6 0.128
LightGBM 10.5 0.826 0.884 7.41 18.5 40.7 55.6 0.128
TabPFN-v2 10.9 0.834 0.923 7.41 22.2 25.9 48.1 0.129
MLP-PLR 11.2 0.827 0.890 0.0 0.0 18.5 40.7 0.131
XGBoost 12.1 0.824 0.875 3.70 18.5 33.3 51.9 0.129
FTT 12.4 0.829 0.900 0.0 11.1 18.5 29.6 0.135
MLP 12.7 0.833 0.915 0.0 7.41 14.8 29.6 0.130
ResNet 13.3 0.834 0.920 0.0 3.70 7.41 29.6 0.130
DCN2 13.4 0.832 0.919 0.0 3.70 3.70 22.2 0.130
PTARL 13.4 0.831 0.911 0.0 3.70 7.41 14.8 0.131
AutoInt 13.7 0.831 0.911 0.0 0.0 3.70 11.1 0.130
Excelformer 14.7 0.821 0.858 3.70 7.41 7.41 11.1 0.143
SNN 14.8 0.831 0.912 0.0 3.70 7.41 18.5 0.131
Laplace KRR 17.0 0.826 0.879 3.70 11.1 14.8 14.8 0.138
Danets 17.4 0.827 0.887 0.0 0.0 0.0 0.0 0.133
TabCaps 18.4 0.822 0.862 0.0 3.70 7.41 7.41 0.136
TabTransformer 19.2 0.816 0.796 0.0 3.70 7.41 18.5 0.160
RandomForest 19.6 0.806 0.772 0.0 3.70 7.41 18.5 0.144
Node 20.1 0.793 0.724 3.70 3.70 3.70 7.41 0.151
LogReg 20.5 0.807 0.778 0.0 3.70 18.5 25.9 0.159
Tangos 20.7 0.811 0.798 0.0 0.0 3.70 3.70 0.139
SVM 21.8 0.802 0.754 0.0 0.0 3.70 18.5 0.162
GrowNet 22.7 0.804 0.771 0.0 0.0 0.0 3.70 0.160
TabNet 23.2 0.814 0.818 0.0 0.0 0.0 0.0 0.142
kNN 24.3 0.787 0.673 0.0 0.0 0.0 7.41 0.166
SwitchTab 27.3 0.786 0.670 0.0 0.0 0.0 0.0 0.163
NaiveBayes 30.2 0.678 0.154 0.0 0.0 0.0 0.0 0.283
NCM 31.3 0.679 0.174 0.0 0.0 0.0 0.0 0.281
Table B.4: Full TALENT Binary classification (>10,000>10,000 samples) results. Rank is the average rank among the ordered methods over all datasets. (Normalized) Score is the metric we use to compare methods in Figure 3. In this case score is the mean classification accuracy (1 minus the error in Figure 3). Top-X (%) is the percentage of datasets for which that method is in the top X ranks. The final column is the shifted geometric mean error ($\mathrm{SGM}_{\varepsilon}$).
Method Rank Score Norm. Score Top-1 (%) Top-3 (%) Top-5 (%) Top-8 (%) $\mathrm{SGM}_{\varepsilon}$
TabPFN-v2 5.02 0.864 0.947 26.9 59.1 72.0 80.6 0.104
xRFM 8.75 0.856 0.893 7.53 29.0 49.5 63.4 0.111
ModernNCA 10.3 0.856 0.886 11.8 22.6 31.2 50.5 0.111
CatBoost 10.3 0.845 0.867 4.30 23.7 34.4 54.8 0.119
LightGBM 10.4 0.845 0.869 7.53 18.3 35.5 54.8 0.119
TabR 10.5 0.851 0.873 4.30 17.2 30.1 43.0 0.116
XGBoost 11.4 0.844 0.851 5.38 18.3 35.5 49.5 0.120
RealMLP 11.7 0.850 0.864 1.08 7.53 18.3 35.5 0.116
FTT 13.3 0.836 0.813 1.08 6.45 17.2 32.3 0.124
MLP-PLR 14.2 0.838 0.817 0.0 3.23 11.8 29.0 0.124
RandomForest 14.3 0.836 0.824 3.23 10.8 19.4 33.3 0.129
AutoInt 15.2 0.834 0.803 0.0 3.23 5.38 16.1 0.129
Tangos 15.4 0.834 0.800 0.0 3.23 7.53 16.1 0.131
DCN2 15.5 0.839 0.816 1.08 2.15 9.68 21.5 0.127
Laplace KRR 15.7 0.836 0.801 1.08 16.1 20.4 26.9 0.133
MLP 16.0 0.836 0.797 1.08 3.23 8.60 17.2 0.132
ResNet 16.1 0.839 0.811 1.08 3.23 8.60 19.4 0.129
TabCaps 16.2 0.835 0.804 1.08 3.23 5.38 14.0 0.130
Excelformer 16.6 0.830 0.778 0.0 1.08 7.53 16.1 0.131
SNN 16.6 0.836 0.800 1.08 3.23 4.30 7.53 0.129
Node 17.0 0.824 0.757 0.0 5.38 15.1 25.8 0.137
PTARL 17.0 0.830 0.780 0.0 1.08 3.23 8.60 0.134
Danets 18.0 0.829 0.759 0.0 3.23 6.45 14.0 0.137
TabTransformer 19.7 0.820 0.715 1.08 1.08 6.45 12.9 0.148
LogReg 20.1 0.819 0.723 3.23 9.68 12.9 17.2 0.146
kNN 20.8 0.811 0.677 5.38 10.8 14.0 18.3 0.158
SVM 21.6 0.816 0.698 2.15 3.23 6.45 10.8 0.150
GrowNet 22.2 0.817 0.714 0.0 0.0 0.0 2.15 0.151
SwitchTab 24.0 0.809 0.675 0.0 0.0 0.0 0.0 0.156
TabNet 25.8 0.802 0.633 0.0 0.0 0.0 1.08 0.157
NaiveBayes 28.6 0.695 0.281 1.08 2.15 3.23 4.30 0.253
NCM 29.9 0.723 0.230 0.0 0.0 0.0 2.15 0.257
Table B.5: Full TALENT Binary classification ($\leq 10,000$ samples) results. Rank is the average rank among the ordered methods over all datasets. (Normalized) Score is the metric we use to compare methods in Figure 3. In this case score is the mean classification accuracy (1 minus the error in Figure 3). Top-X (%) is the percentage of datasets for which that method is in the top X ranks. The final column is the shifted geometric mean error ($\mathrm{SGM}_{\varepsilon}$).
Dataset xRFM XGB CatBoost LGBM RealMLP MLP-PLR MLP-RTDL ResNet-RTDL
Airlines_DepDelay_10M 0.9818 0.9813 0.9796 0.9798 0.9786 0.9795 0.9824 0.9818
Allstate_Claims_Severity 0.6489 0.6547 0.6510 0.6530 0.6495 0.6537 0.6557 0.6537
black_friday 0.6881 0.6807 0.6792 0.6787 0.6859 0.6862 0.6929 0.6892
Buzzinsocialmedia_Twitter 0.2080 0.2134 0.3147 0.2789 0.2566 0.2553 0.2840 0.2906
nyc-taxi-green-dec-2016 0.5834 0.6649 0.6567 0.6489 0.6142 0.6523 0.6657 0.6365
wave_energy 0.0020 0.0918 0.0499 0.0821 0.0024 0.0073 0.0254 0.0434
Yolanda 0.7816 0.8012 0.8094 0.7970 0.7869 0.7897 0.7927 0.7856
Table B.6: Meta-test regression datasets with more than 70,000 samples. Error reported is nRMSE averaged over five train/test splits using pytabkit.
Dataset xRFM XGB CatBoost LGBM RealMLP MLP-PLR MLP-RTDL ResNet-RTDL
airlines 0.3341 0.3292 0.3315 0.3287 0.3344 0.3342 0.3342 0.3339
covertype 0.0257 0.0420 0.0612 0.0333 0.0280 0.0364 0.0404 0.0385
Higgs 0.2640 0.2576 0.2576 0.2549 0.2473 0.2528 0.2515 0.2435
jannis 0.2701 0.2799 0.2816 0.2779 0.2702 0.2764 0.2865 0.2795
MiniBooNE 0.0539 0.0529 0.0538 0.0525 0.0484 0.0504 0.0503 0.0488
numerai28.6 0.4778 0.4812 0.4790 0.4782 0.4800 0.4775 0.4771 0.4800
porto-seguro 0.0380 0.0380 0.0381 0.0381 0.0380 0.0380 0.0380 0.0380
dionis 0.0926 0.1219 0.1041 0.1076 0.0887 0.1257 0.1110 0.0907
Fashion-MNIST 0.0889 0.0928 0.0969 0.0895 0.0913 0.1064 0.1041 0.1011
kick 0.0972 0.0965 0.0956 0.0964 0.0976 0.0978 0.0979 0.0970
Table B.7: Meta-test classification datasets with greater than 70,000 samples. Error reported is classification error averaged over five train/test splits using pytabkit.
Model Elo ($\uparrow$) Norm. Avg. Harm. #wins ($\uparrow$) Improva- Train time Predict time
score ($\uparrow$) rank ($\downarrow$) mean bility ($\downarrow$) per 1K [s] per 1K [s]
rank ($\downarrow$)
AutoGluon 1.3 (4h) 1779 0.673 5.2 3.2 1 3.6% 1734.20 7.06
RealMLP (T+E) 1721 0.660 6.6 5.4 0 3.1% 6860.54 7.68
ModernNCA (T+E) 1626 0.622 9.2 2.6 3 5.4% 3811.43 7.58
LightGBM (T+E) 1602 0.478 10.1 7.5 0 6.0% 686.46 5.48
TabDPT (D) 1575 0.515 10.9 3.8 2 3.9% 16.97 8.70
xRFM (T+E) 1563 0.492 11.5 5.3 1 5.1% 365.57 0.72
CatBoost (T+E) 1558 0.435 11.8 8.9 0 5.7% 2895.38 1.32
TabM (T+E) 1529 0.439 12.8 8.5 0 4.3% 4228.53 1.19
CatBoost (T) $15$ 0.406 13.3 6.8 0 5.9% 2895.38 0.07
xRFM (T) 1504 0.406 13.8 9.7 0 5.7% 365.57 0.09
LightGBM (T) 1498 0.354 13.9 10.2 0 6.8% 686.46 0.74
XGBoost (T+E) 1470 0.328 15.2 13.9 0 6.7% 848.99 2.38
XGBoost (T) 1470 0.325 15.2 13.3 0 6.7% 848.99 0.47
ModernNCA (D) 1457 0.323 15.6 11.0 0 8.1% 16.07 0.29
TabM (T) 1453 0.311 16.0 12.6 0 5.2% 4228.53 0.13
RealMLP (T) 1439 0.306 16.6 13.3 0 5.9% 6860.54 0.32
CatBoost (D) 1419 0.278 17.4 12.0 0 7.9% 8.35 0.09
ModernNCA (T) 1411 0.347 17.6 6.7 0 7.9% 3811.43 0.45
TabPFNv2 (T+E) 1377 0.410 19.1 2.7 4 4.9% 3805.62 10.41
TabM (D) 1361 0.232 20.1 16.5 0 6.9% 13.90 0.12
TabPFNv2 (T) 1314 0.281 22.1 6.9 0 6.4% 3805.62 0.26
TorchMLP (T+E) 1301 0.141 22.7 19.1 0 8.1% 4452.11 0.85
ExtraTrees (T+E) 1297 0.169 23.1 16.9 0 11.1% 161.73 0.78
RealMLP (D) 1274 0.085 24.0 21.6 0 8.2% 23.30 1.44
ExtraTrees (T) 1273 0.174 24.0 16.7 0 11.4% 161.73 0.12
TabPFNv2 (D) 1264 0.262 24.3 7.7 0 7.6% 2.78 0.32
TorchMLP (T) 1245 0.124 25.2 20.8 0 8.7% 4452.11 0.09
xRFM (D) 1236 0.149 25.8 8.6 1 11.7% 1.65 0.08
LightGBM (D) 1231 0.069 25.9 24.5 0 9.7% 2.03 0.30
XGBoost (D) 1195 0.104 27.4 24.5 0 10.4% 2.15 0.18
RandomForest (T+E) 1193 0.056 27.5 25.5 0 12.1% 526.17 0.77
RandomForest (T) 1142 0.046 29.9 27.5 0 12.8% 526.17 0.12
EBM (T+E) 1133 0.122 30.0 16.4 0 14.7% 2124.78 0.12
ExtraTrees (D) 1117 0.056 30.6 27.7 0 13.0% 0.42 0.06
FastaiMLP (T+E) 1086 0.005 31.8 30.6 0 13.1% 527.21 2.83
EBM (T) 1076 0.131 32.2 9.4 1 15.3% 2124.78 0.01
TorchMLP (D) 1075 0.015 32.4 30.3 0 12.9% 20.50 0.08
FastaiMLP (T) 1046 0.000 33.3 32.3 0 13.6% 527.21 0.31
EBM (D) 1009 0.071 34.7 30.3 0 16.0% 7.25 0.04
TabICL (D) 1006 0.000 34.8 34.2 0 14.2% 0.63 0.06
RandomForest (D) 1000 0.000 34.8 34.2 0 14.2% 0.63 0.06
FastaiMLP (D) 916 0.000 37.2 36.5 0 17.8% 3.08 0.29
KNN (T+E) 533 0.000 43.8 43.6 0 37.3% 2.25 0.15
Linear (T+E) 485 0.000 44.3 44.2 0 35.5% 46.50 0.14
KNN (T) 434 0.000 44.9 44.7 0 38.0% 2.25 0.03
Linear (T) 426 0.000 44.9 44.8 0 35.7% 46.50 0.04
Linear (D) 312 0.000 46.0 46.0 0 38.1% 1.16 0.08
KNN (D) 263 0.000 46.5 46.2 0 41.6% 0.04 0.02
Table B.8: Full regression results on TabArena-Lite.
Model Elo ($\uparrow$) Norm. Avg. Harm. #wins ($\uparrow$) Improva- Train time Predict time
score ($\uparrow$) rank ($\downarrow$) mean bility ($\downarrow$) per 1K [s] per 1K [s]
rank ($\downarrow$)
AutoGluon 1.3 (4h) 1521 0.586 9.0 3.0 2 9.7% 4917.81 4.05
CatBoost (T) 1441 0.439 12.5 9.6 0 11.9% 3307.58 0.14
CatBoost (T+E) 1441 0.446 12.4 9.4 0 11.1% 3307.58 1.18
TabPFNv2 (T+E) 1419 0.518 13.5 4.2 1 13.8% 2584.13 12.37
LightGBM (T+E) 1418 0.393 13.2 7.5 0 13.6% 1280.01 4.08
LightGBM (T) 1405 0.381 13.9 6.9 0 13.6% 1280.01 1.05
TabM (T+E) 1387 0.405 14.8 9.7 0 15.7% 5568.31 1.78
XGBoost (T+E) 1377 0.310 15.2 13.3 0 14.0% 2029.77 4.11
TabM (T) 1373 0.401 15.5 9.1 0 16.0% 5568.31 0.37
XGBoost (T) 1373 0.317 15.6 13.1 0 13.7% 2029.77 1.04
RealMLP (T+E) 1359 0.337 16.2 5.3 1 16.8% 6866.35 10.40
TabPFNv2 (D) 1358 0.361 16.4 4.7 1 12.5% 5.48 0.35
xRFM (T+E) 1357 0.369 16.2 7.0 0 14.8% 515.01 1.68
TabPFNv2 (T) 1346 0.366 17.0 6.5 0 14.9% 2584.13 0.41
ModernNCA (T+E) 1345 0.302 16.9 10.4 0 15.3% 6684.65 9.59
ModernNCA (T) 1332 0.248 17.8 12.4 0 15.6% 6684.65 0.75
RealMLP (T) 1304 0.253 19.1 15.5 0 18.4% 6866.35 0.92
CatBoost (D) 1296 0.201 19.2 16.9 0 16.1% 43.10 0.25
TabICL (D) 1293 0.325 19.9 4.5 1 20.5% 11.51 1.95
TabM (D) 1293 0.284 19.9 9.4 0 19.7% 17.09 0.15
ExtraTrees (T+E) 1286 0.289 20.2 11.0 0 17.7% 728.32 2.44
xRFM (T) 1283 0.273 20.4 5.2 1 16.4% 515.01 0.20
ExtraTrees (T) 1265 0.262 21.2 11.2 0 17.3% 728.32 0.36
RandomForest (T+E) 1260 0.321 20.9 6.1 0 15.9% 729.17 1.83
TorchMLP (T+E) 1260 0.188 21.2 17.5 0 19.6% 3646.83 2.16
TorchMLP (T) 1218 0.156 23.9 17.8 0 21.6% 3646.83 0.19
RandomForest (T) 1190 0.241 24.9 12.1 0 17.6% 729.17 0.33
XGBoost (D) 1188 0.102 24.9 21.5 0 18.1% 4.93 0.59
TabDPT (D) 1186 0.259 25.2 4.7 1 23.3% 33.52 20.75
FastaiMLP (T+E) 1182 0.205 25.2 12.1 0 24.2% 2721.87 12.59
LightGBM (D) 1180 0.158 25.4 19.6 0 21.3% 5.12 0.44
EBM (T+E) 1168 0.166 26.0 17.1 0 24.6% 1471.12 0.27
EBM (T) 1144 0.160 27.4 19.0 0 24.9% 1471.12 0.03
xRFM (D) 1129 0.094 28.4 22.1 0 23.9% 2.22 0.20
ModernNCA (D) 1110 0.051 29.1 26.8 0 25.9% 17.24 0.57
FastaiMLP (T) 1100 0.104 29.6 20.3 0 25.8% 2721.87 1.08
RealMLP (D) 1049 0.071 32.0 24.5 0 26.3% 26.02 4.18
EBM (D) 1046 0.095 32.0 24.3 0 27.7% 6.16 0.08
RandomForest (D) 1000 0.000 33.9 32.0 0 33.3% 0.74 0.15
TorchMLP (D) 972 0.004 35.0 33.6 0 28.5% 14.37 0.36
FastaiMLP (D) 952 0.038 35.9 31.9 0 33.4% 8.37 0.66
ExtraTrees (D) 862 0.013 38.6 35.1 0 39.1% 0.76 0.15
Linear (T+E) 752 0.000 41.8 41.5 0 45.6% 170.51 0.20
Linear (T) 718 0.000 42.3 41.9 0 45.8% 170.51 0.13
KNN (T+E) 714 0.000 42.4 41.0 0 52.5% 2.99 0.17
Linear (D) 689 0.000 42.9 42.7 0 46.9% 3.89 0.16
KNN (D) 544 0.000 45.4 45.1 0 69.5% 0.33 0.05
KNN (T) 541 0.000 45.6 45.5 0 58.3% 2.99 0.06
Table B.9: Full multiclass results on TabArena-Lite.
Model Elo ($\uparrow$) Norm. Avg. Harm. #wins ($\uparrow$) Improva- Train time Predict time
score ($\uparrow$) rank ($\downarrow$) mean bility ($\downarrow$) per 1K [s] per 1K [s]
rank ($\downarrow$)
AutoGluon 1.3 (4h) 1779 0.673 5.2 3.2 1 3.6% 1734.20 7.06
RealMLP (T+E) 1721 0.660 6.6 5.4 0 3.1% 6860.54 7.68
ModernNCA (T+E) 1626 0.622 9.2 2.6 3 5.4% 3811.43 7.58
LightGBM (T+E) 1602 0.478 10.1 7.5 0 6.0% 686.46 5.48
TabDPT (D) 1575 0.515 10.9 3.8 2 3.9% 16.97 8.70
xRFM (T+E) 1563 0.492 11.5 5.3 1 5.1% 365.57 0.72
CatBoost (T+E) 1558 0.435 11.8 8.9 0 5.7% 2895.38 1.32
TabM (T+E) 1529 0.439 12.8 8.5 0 4.3% 4228.53 1.19
CatBoost (T) $15$ 0.406 13.3 6.8 0 5.9% 2895.38 0.07
xRFM (T) 1504 0.406 13.8 9.7 0 5.7% 365.57 0.09
LightGBM (T) 1498 0.354 13.9 10.2 0 6.8% 686.46 0.74
XGBoost (T+E) 1470 0.328 15.2 13.9 0 6.7% 848.99 2.38
XGBoost (T) 1470 0.325 15.2 13.3 0 6.7% 848.99 0.47
ModernNCA (D) 1457 0.323 15.6 11.0 0 8.1% 16.07 0.29
TabM (T) 1453 0.311 16.0 12.6 0 5.2% 4228.53 0.13
RealMLP (T) 1439 0.306 16.6 13.3 0 5.9% 6860.54 0.32
CatBoost (D) 1419 0.278 17.4 12.0 0 7.9% 8.35 0.09
ModernNCA (T) 1411 0.347 17.6 6.7 0 7.9% 3811.43 0.45
TabPFNv2 (T+E) 1377 0.410 19.1 2.7 4 4.9% 3805.62 10.41
TabM (D) 1361 0.232 20.1 16.5 0 6.9% 13.90 0.12
TabPFNv2 (T) 1314 0.281 22.1 6.9 0 6.4% 3805.62 0.26
TorchMLP (T+E) 1301 0.141 22.7 19.1 0 8.1% 4452.11 0.85
ExtraTrees (T+E) 1297 0.169 23.1 16.9 0 11.1% 161.73 0.78
RealMLP (D) 1274 0.085 24.0 21.6 0 8.2% 23.30 1.44
ExtraTrees (T) 1273 0.174 24.0 16.7 0 11.4% 161.73 0.12
TabPFNv2 (D) 1264 0.262 24.3 7.7 0 7.6% 2.78 0.32
TorchMLP (T) 1245 0.124 25.2 20.8 0 8.7% 4452.11 0.09
xRFM (D) 1236 0.149 25.8 8.6 1 11.7% 1.65 0.08
LightGBM (D) 1231 0.069 25.9 24.5 0 9.7% 2.03 0.30
XGBoost (D) 1195 0.104 27.4 24.5 0 10.4% 2.15 0.18
RandomForest (T+E) 1193 0.056 27.5 25.5 0 12.1% 526.17 0.77
RandomForest (T) 1142 0.046 29.9 27.5 0 12.8% 526.17 0.12
EBM (T+E) 1133 0.122 30.0 16.4 0 14.7% 2124.78 0.12
ExtraTrees (D) 1117 0.056 30.6 27.7 0 13.0% 0.42 0.06
FastaiMLP (T+E) 1086 0.005 31.8 30.6 0 13.1% 527.21 2.83
EBM (T) 1076 0.131 32.2 9.4 1 15.3% 2124.78 0.01
TorchMLP (D) 1075 0.015 32.4 30.3 0 12.9% 20.50 0.08
FastaiMLP (T) 1046 0.000 33.3 32.3 0 13.6% 527.21 0.31
EBM (D) 1009 0.071 34.7 30.3 0 16.0% 7.25 0.04
TabICL (D) 1006 0.000 34.8 34.2 0 14.2% 0.63 0.06
RandomForest (D) 1000 0.000 34.8 34.2 0 14.2% 0.63 0.06
FastaiMLP (D) 916 0.000 37.2 36.5 0 17.8% 3.08 0.29
KNN (T+E) 533 0.000 43.8 43.6 0 37.3% 2.25 0.15
Linear (T+E) 485 0.000 44.3 44.2 0 35.5% 46.50 0.14
KNN (T) 434 0.000 44.9 44.7 0 38.0% 2.25 0.03
Linear (T) 426 0.000 44.9 44.8 0 35.7% 46.50 0.04
Linear (D) 312 0.000 46.0 46.0 0 38.1% 1.16 0.08
KNN (D) 263 0.000 46.5 46.2 0 41.6% 0.04 0.02
Table B.10: Full binary classification results on TabArena.
Figure B.1: Runtime comparison as a function of the number of training samples on the covertype dataset. Here, L1 kernel refers to the $L_{p}^{p}$ kernel.